# Synthetic Datasets for Machine Learning on Spatio-Temporal Graphs using PDEs

### v1

@@ Abstract
Many physical processes can be expressed through partial differential equations (PDEs). Real-world measurements of such processes are often collected at irregularly distributed points in space, which can be effectively represented as graphs; however, there are currently only a few existing datasets. Our work aims to make advancements in the field of PDE-modeling accessible to the temporal graph machine learning community, while addressing the data scarcity problem, by creating and utilizing datasets based on PDEs. In this work, we create and use synthetic datasets based on PDEs to support spatio-temporal graph modeling in machine learning for different applications. More precisely, we showcase three equations to model different types of disasters and hazards in the fields of epidemiology, atmospheric particles, and tsunami waves. Further, we show how such created datasets can be used by benchmarking several machine learning models on the epidemiological dataset. Additionally, we show how pre-training on this dataset can improve model performance on real-world epidemiological data. The presented methods enable others to create datasets and benchmarks customized to individual requirements. The source code for our methodology and the three created datasets can be found on github.com/github-usr-ano/Temporal_Graph_Data_PDEs.
@@ 1 Introduction
Graphs are a powerful tool for describing complex geometric structures in real-world scenarios, such as surfaces, volumes, or sensor networks. Recently, there have been notable developments in specifically spatio-temporal graph modeling within the field of machine learning (Chen and Eldardiry, 2024). One big obstacle for benchmarking, comparing, and for further developments, is the availability of spatio-temporal data. Only few datasets for spatio-temporal graph machine learning can be found in literature, such as traffic data from the California PeMS system (Li et al., 2018; Yu et al., 2018; Cini et al., 2023) in different settings, or COVID-19 data Gao and Ribeiro (2022). For some benchmarking applications Huang et al. (2023); Gastinger et al. (2024) offer different datasets of interest. Although the existent datasets are valuable for specific tasks, they have limitations in quality (high noise), scope (few samples), application (topic), accessibility (unpublished), and adaptability (high individual pre-processing effort). Thus, we identify a lack of larger high-quality datasets for different tasks. These limitations hinder the development and comprehensive benchmarking of spatio-temporal graph architectures for specific applications and diverse tasks. Therefore, large and more diverse high-quality datasets that are freely accessible are crucial.
@@ 2 Temporal PDEs to Simulate Spatio-Temporal Movement
We solve the equations iteratively in the time domain and solve the occurring PDE for each time-step with the FEM, this order is called Rothe method. We therefore first discretize alongside the temporal dimension with an Euler scheme, i.e. approximate the time-derivative with a difference quotient. We use the Crank-Nicolson method, which is a mixture of the implicit and explicit Euler scheme for time-stepping. With $\theta=0.5$, a small temporal Euler stepsize $h$, and $\nu=\nu(x,t)$ and $\nu_{+}=\nu(x,t+h)$ the approximation has the form
@@ 3 Dataset Creation
As an underlying domain for both the epidemiological PDE and the advection-diffusion PDE, we use the shape of Germany. We evaluate the solutions of the PDEs on a set of points that have the coordinates of administrative regions in Germany: $400$ NUTS-3 regions. The intuition is the reporting of COVID-19 cases in the same spatial granularity. We additionally manually create an adjacency list of neighboring regions and weight them with the inverse of the distance of the centers of the regions, More details on this can be found in the appendix A. We ultimately have created a graph
@@
with $400=|V|$ nodes, $2088=|E|$ edges and edge features $X_{i}\in\mathbb{R}^{d}$ for every time-step $i$. The mesh for the FEM computation and the resulting graph can be seen in Figure 1. Note, that through the use of the FEM our method is not limited to fixed graphs. The values of $X_{i}$ are created by evaluating the solution $\nu(x,t)$ on the points of $V$. No changes to the PDE solver needs to be made, to evaluate the solution on different locations every time-step. However, to correspond to real world data, we chose a fix $V$. On our GitHub we provide further explanations on how to increase and change the size of the graph $V$ with minimal effort. Note that no recomputation of $\nu(x,t)$ is needed to create larger graphs, and the evaluation of $\nu$ can easily be parallelized, thus our method is suitable to create also very large graphs.
@@
We set the parameter $\alpha=0.22$, the time-weighting parameter $\Theta=0.5$ effectively resulting in a Crank-Nicolson scheme, the Euler stepsize $h_{k}=1$, and a Newton stepsize $\eta=0.3$. The underlying domain $\Omega$ can be seen in Figure 1.
We generated $25$ scenarios, each with a length of $364$ time-steps. For every simulation we varied the parameters $r,D$, they can be found in the supplementary material in A and Table 3, leading to a dataset of shape $[9100,400,2]$ for [time-steps, nodes, features] in $5$:$27$ hours (wall clock) of computation. Note, that two features exist due to $2=d=|\{S,I\}|$.
@@
where $\Omega_{s(x,t)}$ is a rectangle within the domain. The choice of $h,T,S(t),\beta(x,t),\Omega_{s(x,t)}$ and further details can be found in the supplementary material A.2 Table 4, resulting in a dataset of shape $[4320,400,1]$, for [time-steps, nodes, features] in 00:$22$h. A spatio-temporal visualization can be found in 4.
@@
For the wave equation, we simulated one consecutive simulation in which we simulated two tsunami waves from different starting points with an initial amplitude of $1$. We set $T=2700000$ and $h=100000/64$, resulting in a dataset of shape $[1858,325,1]$, for [time-steps, nodes, features] in 00:$20$h. Details can be found in the appendix A.3.
@@ 4.2 Epidemiological Benchmarking
We will proceed with the epidemiological dataset which we split $80/20$ along the time axis into a train and a test dataset, and omit any information about the Susceptible $S$, as this reflects real-world data. On this data, we define the following tasks for machine learning models.
@@
Table 1: Overview of Performance compared to different tasks. We take the RMSEs over all samples and all forecasted time-steps from the test dataset. The $\pm$ indicates the standard deviation out of 3 training runs with random initialization respectively. We scaled all values up by the factor $100$ for readability.
@@
Repetition $2.27$ $2.74$ $3.30$ $2.74$ $3.30$
MP-PDE $1.09\pm 0.15$ $2.2\pm 0.2$ $2.15\pm 0.1$ $1.37\pm 0.05$ $\mathbf{1.34\pm 0.18}$
RNN-GNN-Fusion $\mathbf{0.97\pm 0.10}$ $\mathbf{1.3\pm 0.1}$ $\mathbf{1.80\pm 0.1}$ $\mathbf{1.33\pm 0.06}$ $1.48\pm 0.11$
RNN $1.67\pm 0.47$ $2.3\pm 0.3$ $2.81\pm 0.2$ $1.76\pm 0.08$ $2.98\pm 0.20$
GraphEncoding $3.51\pm 0.67$ $3.5\pm 0.7$ $4.10\pm 0.5$ $3.81\pm 0.66$ $5.36\pm 1.12$
TST $1.14\pm 0.03$ $2.6\pm 0.1$ $2.5\pm 0.02$ $2.98\pm 0.68$ $2.72\pm 0.02$
@@ 4.3 Transfer Learning to Real-World Data
While the benchmarking of different model architectures already provides valuable insights, we aim to further illustrate the utility of synthetic data in developing foundational forecasting models for epidemiology or pre-training models on tasks on which data is sparse. We, therefore, seek to demonstrate whether pre-training on our provided synthetic data can lead to improved performance on real-world tasks. To carry out this experiment we train and evaluate the models from the previous section on real-world epidemiological data from respectively German COVID-19, German Influenza, and Brazilian COVID-19. The prediction task is again a 14-day forecast based on 14 days of input. Simultaneously, we pre-train the same models first on the synthetic dataset based on the SI-diffusion equation, and only then shortly retrain (fine-tune) them on the respective real-world data. We then compare the difference in performance. Note, that the Brazilian data requires an additional knowledge transfer onto an unseen domain (i.e. Graph). Details on the real-world datasets can be found in section A.4.
The outcomes of the transfer-learning experiments were highly favorable for pre-training on our dataset, underscoring the effectiveness and possible impact of our proposed method of data synthetization on real-world data and problems. The results can be found in Table 2. A graphic representation, containing the actual RMSEs and comparing the models against each other can be found in the appendix, see fig. 13.
Table 2: Change in validation-loss of the models in the transfer-learning tasks in percent. A negative number indicates improved performance through out method, i.e. by pre-training on synthetic data as opposed to training solely on the actual data.
@@ 5 Conclusions
Our datasets enables others to build, test, and compare new models for three different applications on large amounts high-quality data, stemming from a fully controllable simulation. A comparable epidemiological benchmarking with publicly accessible data is unknown to the authors, and especially conducting such a benchmarking with a comparable amount of data was not possible prior to our contribution. We also strongly encourage others to adapt the code and methodology, such as geometry, parameters, scenarios, or PDEs. This not only enables researchers to fulfill individual needs, but also enables others to explore the capacity of machine learning models to simulate specific behaviors of PDEs.
@@
This work was supported by the Federal Ministry for Economic Affairs and Climate Action (BMWK) as grant DAKI-FWS (01MK21009A).
@@ Appendix A Details on: Dataset Creation
The evaluation points to create the Graph $\mathcal{G}$ were created by finding the centers of the German NUTS3 regions. Geographical data regarding the regions are publicly available. 222 mis.bkg.bund.de/trefferanzeige?docuuid=D38F5B40-9209-4DC0-82BC-57FB575D29D7, which is published under the dl-de/by-2-0 license (https://www.govdata.de/dl-de/by-2-0 ) The data was processed with the use of GeoPandas (Jordahl et al., 2020), shapely and SciPy (Virtanen et al., 2020) by finding centers of each region, adjacent regions and their distances. Some code regarding this can be found in a notebook on our GitHub under /additional_resources/point_and_mesh_generation.ipynb. The underlying graph of the wave-equation dataset was created with the same geometry files. The evaluation points were chosen randomly, the adjacencies were created using a Delaunay triangulation.
@@ A.2 Advection-Diffusion equation
The support, necessary to describe the source term $s(x,t)$, $\Omega_{s(x,t)}$, is a rectangle $[a\cdot 10^{6},1.01a\cdot 10^{6}]\times[b\cdot 10^{6},1.01b\cdot 10^{6}]$ on the domain.
@@
Where $I\in\mathbb{R}$ is some interval. Also $\beta(x,t)$, and $a,b$ can be found in Table 4.

### v2

@@ Abstract
Many physical processes can be expressed through partial differential equations (PDEs). Real-world measurements of such processes are often collected at irregularly distributed points in space, which can be effectively represented as graphs; however, there are currently only a few existing datasets. Our work aims to make advancements in the field of PDE-modeling accessible to the temporal graph machine learning community, while addressing the data scarcity problem, by creating and utilizing datasets based on PDEs. In this work, we create and use synthetic datasets based on PDEs to support spatio-temporal graph modeling in machine learning for different applications. More precisely, we showcase three equations to model different types of disasters and hazards in the fields of epidemiology, atmospheric particles, and tsunami waves. Further, we show how such created datasets can be used by benchmarking several machine learning models on the epidemiological dataset. Additionally, we show how pre-training on this dataset can improve model performance on real-world epidemiological data. The presented methods enable others to create datasets and benchmarks customized to individual requirements. The source code for our methodology and the three created datasets can be found on github.com/Jostarndt/Synthetic_Datasets_for_Temporal_Graphs.
@@ 1 Introduction
Graphs are a powerful tool for describing complex geometric structures in real-world scenarios, such as surfaces, volumes, or sensor networks. Recently, there have been notable developments in specifically spatio-temporal graph modeling within the field of machine learning (Chen and Eldardiry, 2024). One big obstacle for benchmarking, comparing, and for further developments, is the availability of spatio-temporal data. Only few datasets for spatio-temporal graph machine learning can be found in literature, such as traffic data from the California PeMS system (Li et al., 2018; Yu et al., 2018; Cini et al., 2023) in different settings, or COVID-19 data Gao and Ribeiro (2022). For some benchmarking applications Huang et al. (2023); Gastinger et al. (2024) offer different datasets of interest, Rozemberczki et al. (2021) offers access to some datasets with varying size, scope and quality. Although the existent datasets are valuable for specific tasks, they have limitations in quality (high noise), scope (few samples), application (topic), accessibility (unpublished), and adaptability (high individual pre-processing effort).
Thus, we identify a lack of larger high-quality datasets for different tasks. These limitations hinder the development and comprehensive benchmarking of spatio-temporal graph architectures for specific applications and diverse tasks. Therefore, large and more diverse high-quality datasets that are freely accessible are crucial.
@@ 2 Temporal PDEs to Simulate Spatio-Temporal Movement
We solve the equations iteratively in the time domain and solve the occurring PDE for each time-step with the FEM, this order is called Rothe method. We therefore first discretize alongside the temporal dimension with an Euler scheme, i.e. approximate the time-derivative with a difference quotient. We use the Crank-Nicolson method, which is a mixture of the implicit and explicit Euler scheme for time-stepping, which is known for good stability and accuracy with still only moderate implementation complexity. With $\theta=0.5$ (Crank-Nicolson), a small temporal Euler stepsize $h$, and $\nu=\nu(x,t)$ and $\nu_{+}=\nu(x,t+h)$ the approximation has the form
@@
Further, we want to mention the most prominent downside of the FEM, which is its poor scaling behavior for high-dimensional problems: the number of support points usually increases exponentially with the dimensionality of the domain (curse of dimensionality). Additionally the matrix $A$ grows quadratically (but is extremely sparse) with the amount of support points. However, fine meshes are generally advantageous for obtaining accurate solutions. To address these computational challenges, various techniques have been developed, including adaptive mesh refinement, algorithms for sparse matrices, matrix-free solvers, and specialized computational tricks for individual PDEs. While these considerations are essential, an in-depth discussion of them falls outside the scope of our paper, as they are thoroughly addressed in textbooks and continue to be an active area of research.
@@ 3 Dataset Creation
As an underlying domain $\Omega$ for both the epidemiological PDE and the advection-diffusion PDE, we use the shape of Germany. We evaluate the solutions $\nu(x,t)$ of the PDEs (spatially) on a set of points that have the coordinates of administrative regions in Germany: $400$ NUTS-3 regions. To be more precise we will use the centers of the regions: $V=(x_{j})_{j=1,...,400}\subset\Omega$. The intuition is the reporting of COVID-19 cases in the same spatial granularity. The FEM allows us the flexible evaluation of $\nu$ everywhere on the domain. The temporal discretization with the Euler method allows the evaluation of $\nu(x,t)$ on all time-points $t=t_{0}+ih=ih$ (where $h$ is the Euler stepsize as introduced above), which we shortly refer to as timestep $i$. Thus, for each time-step $i$ we can evaluate $\nu(x,t)$ at every point $x_{j}$ and assign a feature vector to every such region, leading to a tensor $X_{i}\in\mathbb{R}^{400\times d}$. For every location $x_{j}$ this tensor contains an entry, which is the evaluation of the PDE-solution, i.e. the entry of $X_{i}$ at position $j$ is $\nu(x_{j},ih)\in\mathbb{R}^{d}$. To equip this data with geographical information, we create a set of edges $E$, containing the edge $(x_{i},x_{j})$ (and $(x_{j},x_{i})$ since its undirected), iff the two regions $x_{i},x_{j}$ directly neighbour each other (are adjacent). We further weight these edges with the inverse of the distances of the centers of the regions in metres. More details on this can be found in the appendix A. We thus have created a graph
@@
with $400=|V|$ nodes, $2088=|E|$ edges and edge features. The mesh for the FEM computation and the resulting graph can be seen in Figure 1. Note, that through the use of the FEM our method is not limited to fixed graphs: since the values of $X_{i}$ are created by evaluating the solution $\nu(x,t)$ on the points of $V$, no changes to the PDE solver needs to be made, to evaluate the solution on different locations every time-step, e.g. make $V$ and thus $E$ time-dependent and create a graph $((V_{i},E_{i},X_{i}))_{i=1,...N}$. However, to correspond to real world data, we chose a fix $V$. On our GitHub we provide further explanations on how to increase and change the size of the graph $V$ with minimal effort. Note that no re-computation of $\nu(x,t)$ is needed to create larger graphs, and the evaluation of $\nu$ can easily be parallelized, thus our method is suitable to create also very large graphs.
@@
We set the parameter $\alpha=0.22$, the time-weighting parameter $\Theta=0.5$ effectively resulting in a Crank-Nicolson scheme for accuracy and stability, the Euler stepsize $h_{k}=1$ (we fixed this first and only then determined the time-dependent dynamics), and a Newton stepsize $\eta=0.3$ for stability and convergence. For other PDEs and parameters other choices have to be determined, that cannot be discussed in the scope of this paper, but are subject to the topic of numerical analysis. The underlying domain $\Omega$ can be seen in Figure 1.
We generated $25$ scenarios, each with a length of $364$ time-steps, resembling an infection scenario over approximately a year. The parameters were chosen heuristically according to their epidemiological interpretation (e.g. disease length around 4.5 days). For every simulation we varied the parameters $r,D$, they can be found in the supplementary material in A and Table 3, leading to a dataset of shape $[9100,400,2]$ for [time-steps, nodes, features] in $5$:$27$ hours (wall clock) of computation. Note, that two features exist due to $2=d=|\{S,I\}|$.
@@
where $\Omega_{s(x,t)}$ is a rectangle within the domain. The choice of $h,T,S(t),\beta(x,t),\Omega_{s(x,t)}$ and further details can be found in the supplementary material A.2 Table 4, resulting in a dataset of shape $[4320,400,1]$, for [time-steps, nodes, features] in 00:$22$h. A spatio-temporal visualization can be found in 4. All selected parameters for this PDE were chosen to generate meaningful and interesting patterns in this visualization method with reasonable sampling rates and frequencies for human eyes and possible ML algorithms.
@@
For the wave equation, we simulated one consecutive simulation in which we simulated two tsunami waves from different starting points with an initial amplitude of $1$. We set $T=2700000$ and $h=100000/64$, resulting in a dataset of shape $[1858,325,1]$, for [time-steps, nodes, features] in 00:$20$h. These parameters were again chosen to have reasonable sampling rates on this comparably large domain, and generate visuably confirmable patterns. Details can be found in the appendix A.3.
@@ 4.2 Epidemiological Benchmarking
We will proceed with the epidemiological dataset which we split $76/12/12$ along the time axis into a train, test and validation dataset. The splits are multiples of four, since $4\%$ of the dataset are exactly one wave/infectious scenario since we simulated 25 different scenarios, and wanted to prevent data-leakage across scenarios and evaluate on full waves. Further we omit any information about the Susceptible $S$, as this reflects real-world data. On this data, we define the following tasks for machine learning models.
@@
Table 1: Overview of Performance compared to different tasks. We take the RMSEs over all samples and all forecasted time-steps from the validation dataset. The $\pm$ indicates the standard deviation out of 3 training runs with random initialization respectively. We scaled all values up by the factor $100$ for readability.
@@
Repetition $2.43\pm 0$ $2.82\pm 0$ $3.27\pm 0.01$ $2.82\pm 0$ $3.27\pm 0.01$
MP-PDE $\mathbf{1.08\pm 0.04}$ $2.04\pm 0.1$ $2.21\pm 0.1$ $1.34\pm 0.01$ $\mathbf{1.11\pm 0.09}$
RNN-GNN-Fusion $1.14\pm 0.16$ $\mathbf{1.30\pm 0.04}$ $\mathbf{1.70\pm 0.1}$ $\mathbf{1.28\pm 0.07}$ $1.58\pm 0.16$
RNN $1.77\pm 0.10$ $1.88\pm 0.1$ $2.84\pm 0.1$ $1.71\pm 0.16$ $3.2\pm 0.11$
GraphEncoding $4.40\pm 1.03$ $4.47\pm 1.03$ $4.86\pm 0.84$ $5.67\pm 3.36$ $4.59\pm 0.28$
TST $1.24\pm 0.13$ $2.45\pm 0.1$ $2.58\pm 0.1$ $2.37\pm 0.19$ $2.69\pm 0.04$
@@ 4.3 Transfer Learning to Real-World Data
While the benchmarking of different model architectures already provides valuable insights, we aim to further illustrate the utility of synthetic data in developing foundational forecasting models for epidemiology or pre-training models on tasks on which data is sparse. We, therefore, seek to demonstrate whether pre-training on our provided synthetic data can lead to improved performance on real-world tasks. To carry out this experiment we train and evaluate the models from the previous section on three real-world epidemiological datasets: (1) German COVID-19, (2) German Influenza, and (3) Brazilian COVID-19. The prediction task is again a 14-day forecast based on 14 days of input. Simultaneously, we pre-train the same models first on the synthetic dataset based on the SI-diffusion equation, and only then shortly retrain (fine-tune) them on the respective real-world data. We then compare the difference in performance. Note, that the Brazilian data requires an additional knowledge transfer onto an unseen domain (i.e. Graph). Details on the real-world datasets can be found in section A.4. The intuition behind this experiment is to test, whether the epidemiological principles, known to humans and expressed in the form of PDEs (with only few parameters), can be taught to machine learning models to generalize onto much more complex real-world data. More general, this transfer-learning experiment tests if the knowledge, encoded in the PDEs, can be used in data-driven applications through pre-training and transfer learning.
The outcomes of the transfer-learning experiments were highly favorable for pre-training on our dataset, underscoring the effectiveness and possible impact of our proposed method of data synthetization on real-world data and problems. Even with the very limited parameter space of the epidemiological PDE, the models learned epidemiological principles and rules expressed through the PDE, which led to improved performances in the real-world scenario. This underlines the importance and potential to integrate knowledge from the field of modeling with PDEs, to the field of temporal graph machine learning in general. Further, this shows how our method addresses the data scarcity problem in this field and potentially other fields. The results can be found in Table 2. A graphic representation, containing the actual RMSEs and comparing the models against each other can be found in the appendix, see fig. 13.
Table 2: Change in validation-loss of the models in the transfer-learning tasks in percent. A negative number indicates improved performance through our method, i.e. by pre-training on synthetic data as opposed to training solely on the actual data.
@@
5 Access to code and data
The code is published under the GNU LESSER GENERAL PUBLIC LICENSE v2.1. 222www.gnu.org/licenses/old-licenses/lgpl-2.1.html. The created datasets are published under the CC BY 4.0 license. 333creativecommons.org/licenses/by/4.0/legalcode.en. Both the code and data will be stored on GitHub, which currently allows for indefinite hosting, ensuring long-term availability as long as GitHub remains operational.
@@ 6 Conclusions
Our datasets enables others to build, test, and compare new models for three different applications on large amounts high-quality data, stemming from a fully controllable simulation. We believe this controllability is also advantageous in reducing human biases, since such bias is more likely to stand-out through the compact and explicit representation of PDEs compared to implicit biases in other data-collection processes. However, we want to mention that theoretically our method could still transfer a bias from PDE-based modeling to data-driven applications through a transfer learning task, potentially leading to biased performance even after fine-tuning, although we consider this risk to be very limited.A comparable epidemiological benchmarking with publicly accessible data is unknown to the authors, and especially conducting such a benchmarking with a comparable amount of data was not possible prior to our contribution. We also strongly encourage others to adapt the code and methodology, such as geometry, parameters, scenarios, or PDEs. This not only enables researchers to fulfill individual needs, but also enables others to explore the capacity of machine learning models to simulate specific behaviors of PDEs.
@@
This work was supported by the Federal Ministry for Economic Affairs and Climate Action (BMWK) as grant DAKI-FWS (01MK21009A); the Senate of Berlin and the European Commision’s Digital Europe Programme (DIGITAL) as grant TEF-Health (101100700), the European Union’s Horizon Europe research and innovation program (EU Horizon Europe) as grant MedEWSa (101121192), and the Fraunhofer Internal Programs as grant PREPARE (40-08394).
@@ Appendix A Details on: Dataset Creation
The evaluation points to create the Graph $\mathcal{G}$ were created by finding the centers of the German NUTS3 regions. Geographical data regarding the regions are publicly available. 444 mis.bkg.bund.de/trefferanzeige?docuuid=D38F5B40-9209-4DC0-82BC-57FB575D29D7, which is published under the dl-de/by-2-0 license (https://www.govdata.de/dl-de/by-2-0 ) The data was processed with the use of GeoPandas (Jordahl et al., 2020), shapely and SciPy (Virtanen et al., 2020) by finding centers of each region, adjacent regions and their distances. Some code regarding this can be found in a notebook on our GitHub under /additional_resources/point_and_mesh_generation.ipynb. The underlying graph of the wave-equation dataset was created with the same geometry files. The evaluation points were chosen randomly, the adjacencies were created using a Delaunay triangulation.
@@ A.2 Advection-Diffusion equation
The support $\Omega_{s(x,t)}$, necessary to describe the source term $s(x,t)$, is a rectangle $[a\cdot 10^{6},1.01a\cdot 10^{6}]\times[b\cdot 10^{6},1.01b\cdot 10^{6}]$ on the domain.
@@
Where $I\in\mathbb{R}$ is some interval. Also, $\beta(x,t)$, and $a,b$ can be found in Table 4.